# String islands, discrete theta angles and the 6D $\mathcal{N}=(1,1)$ string landscape

### v1

@@ 1 Introduction
[…]
@@
However, the 6d $\mathcal{N}=(1,1)$ stringy landscape remains incomplete and hence motivating this work [^1] Note that the chiral 6d theory is unique due to anomalies and can realized as type IIB on $K3$.. A way to realize a large class of these theories is using asymmetric orbifolds with chiral actions [^2] By chiral actions we mean orbifold actions acting exclusively on the worldsheet right-movers or left-movers, excluding quasicrystalline actions [13] which where classified in [14]. [15] of either heterotic or type II strings, with the latter sometimes admitting nonperturbative discrete theta angles [12]. For $\mathbb{Z}_{n}$-orbifolds, it was conjectured in [16] that, at the worldsheet level (i.e. up to discrete theta angle), the heterotic strings correspond to families of currents of chiral 2D conformal field theories with central charge 24 [17, 18, 19], while the type II strings correspond to currents of the chiral 2D superconformal field theory with central charge 12 given by 24 real fermions [20]. Here we present an explicit proof of this statement, which extends to non-cyclic and even to non-Abelian orbifolds.
In fact, a classification was undertaken in [16] with the expectation of 47 moduli spaces. From these 23 are cyclic orbifolds of both the type II and heterotic string theories and 1 is the trivial Narain moduli space of the heterotic theory. In particular, there are 10 heterotic orbifolds in this classification. From these, 8 correspond to known constructions, correspond to $n=2,\cdots 8$, while the three remaining theories, which are more subtle, were recently constructed in [21]. On the other hand, the predicted type II strings add up to 13 (5 with discrete theta angle), out of which 8 (3 with discrete theta angle) are known in the literature at the moment — more details on both heterotic and type II constructions are given below. Here we will construct explicitly the five new predicted theories. In doing so, we show that the predictions regarding discrete theta angles in [16] (which lie outside the scope of the worldsheet classification proof above) are indeed correct. Interestingly, they correspond to isolated theories, where the only massless degrees of freedom belong to the gravity multiplet and the only scalar is the dilaton. Such theories are called string islands [13].
@@ 1.1 Summary of results and organization of the paper
    We prove that every type II asymmetric orbifold in 6D $\mathcal{N}=(1,1)$, where the two gravitini come from chiral worldsheet fields (i.e. those considered in [28]) is in one-to-one correspondence to an orbifold of the $E_{8}$ superconformal field theory, leading to a classification of such 6D theories at the perturbative level. See the Proposition in Section 2.2 for the precise statement. The bulk of the proof is rather technical and we relegate it to Appendix A. We also show that this proof extends to the heterotic string, and that its results are in perfect agreement with the recent explicit constructions of [21].
@@
    We present a general treatment of discrete theta angles in 6D $\mathcal{N}=(1,1)$ asymmetric orbifolds, and verify the prediction [16] that those with $n=5,8$ admit nontrivial discrete theta angles of order 5 and 2, respectively.
@@
    We show explicitly that the discrete theta angle is correlated with the presence of extra non-BPS strings extending the lattice of charges for fundamental strings according to the order of the discrete theta angle as in [12]. Moreover we explicitly compute the lattices of electric charges for particles and how they are extended due to extra non-BPS states, in perfect agreement with the predicted charge lattices in [16].
@@ 2.2 Classification
Figure 1: Correspondence between asymmetric orbifolds of the type II string on $T^{4}$ and the 2D superconformal field theory $F_{4}$ comprising 24 fermions and a current specified by the given Kac-Moody algebra. The rank of the type II theory is equal to the rank of the corresponding current minus 4, hence the theories in the lower row have rank 0 and are accordingly referred to as string islands following [25]. The symbol $\hat{A}_{m,n}$ denotes the Kac-Moody algebra for $SU(m+1)$ at level $n$, and likewise, $\hat{B}_{n,m}$, $\hat{C}_{m,n}$ and $\hat{G}_{2,m}$ refer to $SO(2m+1)$, $Sp(m)$ and $G_{2}$.
@@ 2.3 Explicit constructions
    $\mathbb{Z}_{2}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{Z}_{3}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{Z}_{4}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{Z}_{6}\,.$ (2.12)
@@
    $\mathbb{Z}_{5}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{Z}_{8}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{Z}_{10}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{Z}_{12}\,.$ (2.13)
@@ 2.3.1 The four conventional orbifolds
$M:\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \tau\to-\frac{1}{\tau}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \rho\to-\frac{1}{\rho}\,.$ (2.17)
@@ 2.3.2 String islands
Note that the $\mathbb{Z}_{2}$ subgroup generated by $M^{4}$ has twist vector $4\,,\phi_{R}=(1,2)\simeq(1,0)$ in reflecting that $M^{4}=(-1)^{F_{L}}$. The $\mathbb{Z}_{4}$ subgroup generated by $M^{2}$ has in turn a twist vector $2\,\phi_{R}=(1/2,1)$, and so $M^{2}$ is the product of the symmetry defining the $\mathbb{Z}_{4}$ asymmetric orbifold and $(-1)^{F_{L}}$. This is the alternative but equivalent lift of the lattice automorphism reflecting $2A_{1}$ to a worldsheet symmetry [30]. The shift vector can be taken as
@@ 2.4 Duality with quasicrystalline orbifolds
$\Phi\mapsto-\Phi\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ H_{\mu\nu\lambda}\mapsto e^{-2\Phi}\widetilde{H}_{\mu\nu\lambda}\,,$ (2.30)
@@
$\begin{array}[]{|c|c|c|}\hline\cr\text{Islands}&(\phi_{L};\phi_{R})&(\phi_{L,c};\phi_{R,c})\\ \hline\cr\mathbb{Z}_{5}&(2/5,4/5;0,0)&(-1/5,-1/5;3/5,3/5)\\ \mathbb{Z}_{8}&(1/4,1/2;0,0)&(-1/8,-1/8;3/8,3/8)\\ \mathbb{Z}_{10}&(1/5,2/5;0,0)&(-1/10,-1/10;3/10,3/10)\\ \mathbb{Z}_{12}&(1/3,1/2;0,0)&(-1/12,-1/12;5/12,5/12)\\ \hline\cr\end{array}$
@@ 3.1 The AOB background
$T:\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \begin{pmatrix}C_{2}\\ B_{2}\end{pmatrix}\mapsto\begin{pmatrix}C_{2}+B_{2}\\ B_{2}\end{pmatrix}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (-1)^{F_{L}}T:\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \begin{pmatrix}C_{2}\\ B_{2}\end{pmatrix}\mapsto\begin{pmatrix}-C_{2}-B_{2}\\ B_{2}\end{pmatrix}\,.$ (3.1)
It follows that, in this construction, it is the combination $B_{2}+2C_{2}$ on which the orbifold symmetry acts as $-1$, and not $C_{2}$. The coupling of a $(p,q)$ string to the 2-form fields can be rewritten as
@@
We now extend the analysis of [12] to the lattice of electric charges for particles. In the AOB, the holonomy gives rise to states with half-integral winding in the twisted sector, and so the charge lattice is extended from the Narain lattice $\Gamma_{1,1}$ to $\Gamma_{1,1}(\tfrac{1}{2})$. Non-BPS strings have half-integral charge with respect to the B-field, giving rise non-BPS particles with integral winding which have effectively the same charge as the twisted states above. The overall electric charge lattice is hence equivalent to that of the original AOB. Compactifying on an extra circle leads to different lattices, however. There are no states with half-integral winding along the extra circle, but the non-BPS strings wrap it extending the extra $\Gamma_{1,1}$ to $\Gamma_{1,1}(\tfrac{1}{2})$. For a $T^{n}$ compactification down to $9-n$ spacetime dimensions, we have then
$\begin{split}\Gamma_{\text{AOB}}\left.\right|_{C_{0}=0}&=\Gamma_{1,1}(\tfrac{1}{2})\oplus\Gamma_{n,n}\,,\\ \Gamma_{\text{AOB}}\left.\right|_{C_{0}=\tfrac{1}{2}}&=\Gamma_{1,1}(\tfrac{1}{2})\oplus\Gamma_{n,n}(\tfrac{1}{2})\,.\\ \end{split}$ (3.3)
@@
This normalization is more agreeable to a comparison with results coming from the frozen singularity frame. In eight dimensions, for example, the ADE type of a frozen singularity in F-theory on an elliptic K3 is $D_{8}$. The AOB is dual to a configuration with two frozen singularities, and the discrete theta angle is seen in the torsional part of the Mordell-Weil group of the elliptic fibration. With the discrete theta angle turned off, the middle cohomology sublattice corresponding to the pair of frozen singularities is $(D_{8}\oplus D_{8})^{+}$ where the $+$ means adding the vector in the $(v,v)$ conjugacy class of $D_{8}\oplus D_{8}$. Turning on the discrete theta angle, the frozen singularity lattice is instead $D_{8}\oplus D_{8}$. Taking the orthogonal complement of each lattice within $\Gamma_{2,18}$ one obtains respectively $\Gamma_{1,1}\oplus\Gamma_{1,1}(2)$ and $\Gamma_{2,2}(2)$, coinciding with our results up to an overall normalization by a factor of 2 [11]. This procedure of taking orthogonal complements of lattices associated to singularities coincides with the formal construction of rank reduction maps in [16] down to six dimensions, and agrees with the results above.
@@ 3.2 The general case
$\psi_{a}\mapsto S^{\prime}_{a}{}^{b}\psi_{b}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \psi_{\bar{a}}\mapsto(S^{\prime-1})^{T}{}_{\bar{a}}{}^{\bar{b}}\psi_{\bar{b}}\,.$ (3.8)
@@
$T(u):\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C_{i}\mapsto C_{i}+u_{i}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ u_{i}\in\mathbb{Z}\,,$ (3.11)
@@
$gT(u):\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C_{i}\mapsto S^{\prime}_{i}{}^{j}(C_{j}+u_{j})\,.$ (3.12)
@@ 3.2.1 Effect on the fundamental string charge lattice
$T(u):\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C_{i\mu\nu}\mapsto C_{i\mu\nu}+u_{i}B_{\mu\nu}\,,$ (3.16)
@@
$gT(u):\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C_{i\mu\nu}\mapsto S^{\prime}_{i}{}^{j}(C_{j\mu\nu}+u_{j}B_{\mu\nu})\,.$ (3.17)
@@ 8 $A_{1}(-1)\oplus A_{3}(-1)$ $3\,A_{1}(-1)\oplus A_{1}(-2)$
In the twisted sectors, the translation along $S^{1}$ introduces a fractional shift on the winding number $w$, so that $w\in\mathbb{Z}+p/n$, $q=1,2,...,n-1$ in the $g^{p}$-twisted sector. These states too have charges on the lattice $I^{*}$, hence each of the twisted sectors has a charge lattice equal to a translate of $\Gamma_{c,\text{untwisted}}$. Putting together the contributions from all the sectors in the orbifold, we obtain the full charge lattice
@@
$\begin{split}I&=\Gamma_{4,4}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Gamma_{2,2}\oplus A_{2}(-1)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Gamma_{2,2}\oplus 2\,A_{1}(-1)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Gamma_{2,2}\oplus A_{2}(-1)\,,\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ A_{4}(-1)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ A_{1}(-1)\oplus A_{3}(-1)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ A_{4}(-1)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ A_{2}(-1)\oplus 2A_{1}(-1)\,,\end{split}$ (3.27)
@@
$T(u):\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C_{\mu}\mapsto C_{\mu}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C_{1234\mu}\mapsto C_{1234\mu}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C_{ij\mu}\mapsto C_{ij\mu}+u_{i}B_{j\mu}+u_{j}B_{i\mu}\,.$ (3.29)
@@
$gT(u):\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \psi_{c,\mu}\mapsto S(\psi_{c,\mu}+\chi_{\mu})\,.$ (3.30)
@@
$\widetilde{\psi}_{c,\mu}=\psi_{c\mu}+\delta_{\mu}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \delta_{\mu}\equiv-(1-S)^{-1}S\chi_{\mu}\,.$ (3.31)
@@
This effect is computationally equivalent to allowing fractional windings along each of the four compact directions, hence prior to the orbifold projection, the Narain lattice $\Gamma_{4,4}$ is extended to $\Gamma_{4,4}(1/m)$. After projecting projecting out states, then, the overall change in the charge lattice $\Gamma_{c}$ due to the discrete theta angle is a scaling $I^{*}\mapsto I^{*}(1/m)$, and so

### v2

@@ 1 Introduction
$D=9:\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ r=17,9,1\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \,;\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ D=8:\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ r=18,10,2\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \,;\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ D=7:\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ r=19,11,7,5,3,1\,.$
@@
However, the 6d $\mathcal{N}=(1,1)$ stringy landscape remains incomplete, hence motivating this work [^1] Note that the chiral 6d theory is unique due to anomalies and can be realized as type IIB on $K3$.. A way to realize a large class of these theories is using asymmetric orbifolds with chiral actions [^2] By chiral actions we mean orbifold actions acting exclusively on the worldsheet right-movers or left-movers, excluding quasicrystalline actions [20] which were classified in [4]. [30] of either heterotic or type II strings, with the latter sometimes admitting nonperturbative discrete theta angles [28]. For $\mathbb{Z}_{n}$-orbifolds, it was conjectured in [16] that, at the worldsheet level (i.e. up to discrete theta angle), the heterotic strings correspond to families of currents of chiral 2D conformal field theories with central charge 24 [35, 26, 24], while the type II strings correspond to currents of the chiral 2D superconformal field theory with central charge 12 given by 24 real fermions [19]. Here we present an explicit proof of this statement, which extends to non-cyclic and even to non-Abelian orbifolds.
In fact, a classification was undertaken in [16] with the expectation of 47 moduli spaces. From these, 23 are cyclic orbifolds of both the type II and heterotic string theories and 1 is the trivial Narain moduli space of the heterotic theory. In particular, there are 10 heterotic orbifolds in this classification. From these, 8 correspond to known constructions, with $n=2,\cdots 8$, while the three remaining theories, which are more subtle, were recently constructed in [3]. On the other hand, the predicted type II strings add up to 13 (5 with discrete theta angle), out of which 8 (3 with discrete theta angle) are known in the literature at the moment — more details on both heterotic and type II constructions are given below. Here we will construct explicitly the five new predicted theories. In doing so, we show that the predictions regarding discrete theta angles in [16] (which lie outside the scope of the worldsheet classification proof above) are indeed correct. Interestingly, they correspond to isolated theories, where the only massless degrees of freedom belong to the gravity multiplet and the only scalar is the dilaton. Such theories are called string islands [20].
@@ 1.1 Summary of results and organization of the paper
    We prove that every type II asymmetric orbifold in 6D $\mathcal{N}=(1,1)$, where the two gravitini come from chiral worldsheet fields (i.e. those considered in [13]) is in one-to-one correspondence with an orbifold of the $E_{8}$ superconformal field theory, leading to a classification of such 6D theories at the perturbative level. See the Proposition in Section 2.2 for the precise statement. The bulk of the proof is rather technical and we relegate it to Appendix A. We also show that this proof extends to the heterotic string, and that its results are in perfect agreement with the recent explicit constructions of [3].
@@
    We present a general treatment of discrete theta angles in 6D $\mathcal{N}=(1,1)$ asymmetric orbifolds, and verify the prediction in [16] that those with $n=5,8$ admit nontrivial discrete theta angles of order 5 and 2, respectively.
@@
    We show explicitly that the discrete theta angle is correlated with the presence of extra non-BPS strings extending the lattice of charges for fundamental strings according to the order of the discrete theta angle as in [28]. Moreover, we explicitly compute the lattices of electric charges for particles and how they are extended due to extra non-BPS states, in perfect agreement with the predicted charge lattices in [16].
@@ 2.2 Classification
Figure 1: Correspondence between asymmetric orbifolds of the type II string on $T^{4}$ and the 2D superconformal field theory $F_{24}$ comprising 24 fermions and a current specified by the given Kac-Moody algebra. The rank of the type II theory is equal to the rank of the corresponding current minus 4, hence the theories in the lower row have rank 0 and are accordingly referred to as string islands following [11]. The symbol $\hat{A}_{m,n}$ denotes the Kac-Moody algebra for $SU(m+1)$ at level $n$, and likewise, $\hat{B}_{n,m}$, $\hat{C}_{m,n}$ and $\hat{G}_{2,m}$ refer to $SO(2m+1)$, $Sp(m)$ and $G_{2}$.
@@ 2.3 Explicit constructions
[…]
@@
[…]
@@ 2.3.1 The four conventional orbifolds
$M:\penalty 10000\ \penalty 10000\ \penalty 10000\ \tau\to-\frac{1}{\tau}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \rho\to-\frac{1}{\rho}\,.$ (2.17)
@@ 2.3.2 String islands
Note that the $\mathbb{Z}_{2}$ subgroup generated by $M^{4}$ has twist vector $4\,\phi_{L}=(1,2)\simeq(1,0)$ in reflecting that $M^{4}=(-1)^{F_{L}}$. The $\mathbb{Z}_{4}$ subgroup generated by $M^{2}$ has in turn a twist vector $2\,\phi_{R}=(1/2,1)$, and so $M^{2}$ is the product of the symmetry defining the $\mathbb{Z}_{4}$ asymmetric orbifold and $(-1)^{F_{L}}$. This is the alternative but equivalent lift of the lattice automorphism reflecting $2A_{1}$ to a worldsheet symmetry [39]. The shift vector can be taken as
@@ 2.4 Duality with quasicrystalline orbifolds
$\Phi\mapsto-\Phi\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ H_{\mu\nu\lambda}\mapsto e^{-2\Phi}\widetilde{H}_{\mu\nu\lambda}\,$, (2.30)
@@
[…]
@@ 3.1 The AOB background
$T:\penalty 10000\ \penalty 10000\ \penalty 10000\ \begin{pmatrix}C_{2}\\ B_{2}\end{pmatrix}\mapsto\begin{pmatrix}C_{2}+B_{2}\\ B_{2}\end{pmatrix}\penalty 10000\ \penalty 10000\ \penalty 10000\ \Rightarrow\penalty 10000\ \penalty 10000\ \penalty 10000\ (-1)^{F_{L}}T:\penalty 10000\ \penalty 10000\ \penalty 10000\ \begin{pmatrix}C_{2}\\ B_{2}\end{pmatrix}\mapsto\begin{pmatrix}-C_{2}-B_{2}\\ B_{2}\end{pmatrix}\,$. (3.1)
It follows that, in this construction, it is the combination $B_{2}+2C_{2}$ on which the orbifold symmetry acts as 1-1, and not $C_{2}$. The coupling of a $(p,q)$ string to the 2-form fields can be rewritten as
@@
We now extend the analysis of [28] to the lattice of electric charges for particles. In the AOB, the holonomy gives rise to states with half-integral winding in the twisted sector, and so the charge lattice is extended from the Narain lattice $\Gamma_{1,1}$ to $\Gamma_{1,1}(\tfrac{1}{2})$. Non-BPS strings have half-integral charge with respect to the B-field, giving rise non-BPS particles with integral winding which have effectively the same charge as the twisted states above. The overall electric charge lattice is hence equivalent to that of the original AOB. Compactifying on an extra circle leads to different lattices, however. There are no states with half-integral winding along the extra circle, but the non-BPS strings wrap it extending the extra $\Gamma_{1,1}$ to $\Gamma_{1,1}(\tfrac{1}{2})$. For a $T^{n}$ compactification down to 9n9-n spacetime dimensions, we have then
[…]
@@
This normalization is more agreeable to a comparison with results coming from the frozen singularity frame. In eight dimensions, for example, the ADE type of a frozen singularity in F-theory on an elliptic K3 is $D_{8}$. The AOB is dual to a configuration with two frozen singularities, and the discrete theta angle is seen in the torsional part of the Mordell-Weil group of the elliptic fibration. With the discrete theta angle turned off, the middle cohomology sublattice corresponding to the pair of frozen singularities is $(D_{8}\oplus D_{8})^{+}$ where the ++ means adding the vector in the $(v,v)$ conjugacy class of $D_{8}\oplus D_{8}$. Turning on the discrete theta angle, the frozen singularity lattice is instead $D_{8}\oplus D_{8}$. Taking the orthogonal complement of each lattice within $\Gamma_{2,18}$ one obtains respectively $\Gamma_{1,1}\oplus\Gamma_{1,1}(2)$ and $\Gamma_{2,2}(2)$, coinciding with our results up to an overall normalization by a factor of 2 [33]. This procedure of taking orthogonal complements of lattices associated to singularities coincides with the formal construction of rank reduction maps in [16] down to six dimensions, and agrees with the results above.
@@ 3.2 The general case
$\psi_{a}\mapsto S^{\prime}_{a}{}^{b}\psi_{b}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \psi_{\bar{a}}\mapsto(S^{\prime-1})^{T}{}_{\bar{a}}{}^{\bar{b}}\psi_{\bar{b}}\,.$ (3.8)
@@
$T(u):\penalty 10000\ \penalty 10000\ \penalty 10000\ C_{i}\mapsto C_{i}+u_{i}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ u_{i}\in\mathbb{Z}\,$, (3.11)
@@
[…]
@@ 3.2.1 Effect on the fundamental string charge lattice
$T(u):\penalty 10000\ \penalty 10000\ \penalty 10000\ C_{i\mu\nu}\mapsto C_{i\mu\nu}+u_{i}B_{\mu\nu}\,$, (3.16)
@@
[…]
@@ 8 $A_{1}(-1)\oplus A_{3}(-1)$ $3\,A_{1}(-1)\oplus A_{1}(-2)$
In the twisted sectors, the translation along $S^{1}$ introduces a fractional shift on the winding number $w$, so that $w\in\mathbb{Z}+p/n$, q=1,2,,n1q=1,2,...,n-1 in the $g^{p}$-twisted sector. These states too have charges on the lattice $I^{*}$, hence each of the twisted sectors has a charge lattice equal to a translate of $\Gamma_{c,\text{untwisted}}$. Putting together the contributions from all the sectors in the orbifold, we obtain the full charge lattice
@@
[…]
@@
$T(u):\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ C_{\mu}\mapsto C_{\mu}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ C_{1234\mu}\mapsto C_{1234\mu}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ C_{ij\mu}\mapsto C_{ij\mu}+u_{i}B_{j\mu}+u_{j}B_{i\mu}\,.$ (3.29)
@@
$gT(u):\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \psi_{c,\mu}\mapsto S(\psi_{c,\mu}+\chi_{\mu})\,.$ (3.30)
@@
$\widetilde{\psi}_{c,\mu}=\psi_{c\mu}+\delta_{\mu}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \delta_{\mu}\equiv-(1-S)^{-1}S\chi_{\mu}\,.$ (3.31)
@@
This effect is computationally equivalent to allowing fractional windings along each of the four compact directions, hence prior to the orbifold projection, the Narain lattice $\Gamma_{4,4}$ is extended to $\Gamma_{4,4}(1/m)$. After projecting out states, then, the overall change in the charge lattice $\Gamma_{c}$ due to the discrete theta angle is a scaling $I^{*}\mapsto I^{*}(1/m)$, and so